OpenStudy (anonymous):

To grow grass in Mel Smith's yard, a landscaper estimates that he will need 55 pounds of grass seed. Each pound will cost Mel $1.83. The landscaper will also charge Mel 6 hours of labor at $12.50 per hour. If Mel pays early, he will be entitled to a cash discount of $3.50. Assuming Mel takes advantage of the discount, how much does he pay the landscaper?

OpenStudy (anonymous):

55 * 1.83 = 100.65 for the seed. 6 * 12.50 = 75 for the labor 100.65 + 75 - 3.50 = 172.15 after the discount.